# On the singular loci of higher secant varieties of Veronese embeddings

### v2

@@ 1. Introduction
    (i) of Theorem 1 is stronger than (i) of Theorems 2 and 3, since it claims smoothness for ‘every point’ in the $m$-subsecant locus $\Sigma_{k,d}(m)$ outside $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. The ‘general point’ condition of (i) in Theorems 2 and 3 can not be deleted (see Example 27). Except this aspect, (i)-(iii) of Theorem 1 can be also derived as applying Theorem 2 directly to the case $m=1$ (recall that curve case has no defectivity in its higher secants).
@@
    (i) in both Theorem 2 and Theorem 3 mostly corresponds to the identifiable case of a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$, which means that there is a unique $k$-tuple of points $x_{1},\dots,x_{k}\in v_{d}({\mathbb{P}}^{m})$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$. In the viewpoint of secant fiber in the incidence (3) in §2, it means that the fiber $p^{-1}(a)$ under $p$ consists of only one element up to permuting $x_{i}$’s. On the other hand, (ii) of Theorems 2 and 3 corresponds to the case of generic non-identifiability of the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ and to the situation of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq\langle v_{d}({\mathbb{P}}^{m})\rangle$. This non-identifiability can happen if $\dim(p^{-1}(a))>0$, or if #$\#p^{-1}(a)\geq 2$ (modulo permutation) in case of $\dim p^{-1}(a)=0$; the former one corresponds to $(d,m)\notin\{(6,2),(3,5),(4,3)\}$ and the latter $(d,m)\in\{(6,2),(3,5),(4,3)\}$.
@@
    Theorem 3 deals with the case in which the given subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ can possibly be defective due to the famous theorem of Alexander-Hirschowitz [1], whereas Theorem 2 concerns the case of always non-defective $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$’s.
  5. (e)
    Two cases $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$ (i.e., (iv) of Theorems 1 and  2), a posteriori, turn out to be the only two exceptional cases which do not follow this trichotomy pattern; in other words, though $k$ belongs to the range of (ii), $\Sigma_{k,d}(m)$ does not provide non-trivial singularity (see also Remark 24). In [15], we in fact show that if $(k,d,n)=(3,4,2),(4,3,3)$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is a del Pezzo $k$-secant variety, that is, a $k$-secant variety of next-to-minimal degree. In this sense, these two cases also belong to a special class with respect to degrees of higher secant varieties. (For basic definitions and results on such varieties, see [11], [10].)
As an application of our main results for $\Sigma_{4,d}(2)$ and $\Sigma_{4,d}(1)$, we obtain the following result on the singular locus of the fourth secant variety of any Veronese embeddings.
@@
    For $d\geq 7$, $\Sigma_{4,d}(1)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\operatorname{Sm}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$. When $d=6$, $\Sigma_{4,d}(1)\subset\operatorname{Sing}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{4,d}(1)\not\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). When $d\leq 5$, $\Sigma_{4,d}(1)\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
@@
The smallest case for singular locus of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ beyond the classical results for $d=2$ or $n=1$ is $(d,n)=(3,2)$, but there is nothing to check in this case, because $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))$ fills up the ambient space ${\mathbb{P}}^{9}$ (i.e., $\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{2})))=\emptyset$). In next case $(d,n)=(3,3)$, by Theorem 5 (i) and (ii), we have
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{3})))=\sigma_{3}(v_{3}({\mathbb{P}}^{3}))~{}.$
For the case $(d,n)=(3,4)$, if $V$ denotes a $5$-dimensional ${\mathbb{C}}$-vector space with ${\mathbb{P}}^{4}={\mathbb{P}}V$, Theorem 5 tells us that the singular locus of the $4$-th secant of $v_{3}({\mathbb{P}}V)$ in ${\mathbb{P}}^{34}$ is precisely the locus of cubic hypersurfaces in 5 variables which are cones with the vertex dimension $\geq 1$ as follows:
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}V)))=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\Sigma_{4,3}(2;{{\mathbb{P}}V})=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\sigma_{4}(v_{3}({\mathbb{P}}^{2}))\big{\}}=\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\langle\,v_{3}({\mathbb{P}}^{2})\,\rangle$
=$=\big{\{}f\in{\mathbb{P}}S^{3}V~{}|~{}\textrm{$f$ defines a cubichypersurface in ${\mathbb{P}}V$ which is a cone with $\dim\operatorname{Vertex}(f)\geq 1$}\big{\}}~{},$
@@
This kind of a simple description of the singular locus can be attained in few more cases (see Corollary 32 for details).
The paper is structured as follows. In §2, as preparation, we recall some preliminaries on $k$-th secant varieties and corresponding incidences, and reveal some geometric properties of subsecant varieties in higher secants of Veronese varieties, which are crucial for the proof of main theorems. In §3, as an illustration of the whole picture and our main ideas, we treat the case $m=1$ and prove Theorem 1. In §4 we deal with the general case (i.e., $m\geq 2$) and prove Theorem 2 and Theorem 3 as generalizing the ideas used in the previous section. We’d like to remark that this can be done because the dimension theorem [1] and general identifiability question [9, 12] was settled for the case of Veronese varieties. In §5 we focus on the singular locus of $4$-th secant variety and prove Theorem 5 by dividing the case into two parts; ‘full-secant locus points (i.e., $m=3$)’ treated in Theorem 29 via Young flattening and conormal space computation and ‘the subsecant locus’ by Corollary 30. Finally, we make some generalization and remarks on the materials for further development in §6.
@@ Acknowledgements
The authors would like to give thank to Korea Institute for Advanced Study (KIAS) for inviting them and giving a chance to start the project. We are grateful to Giorgio Ottaviani for a helpful comment on Example 6 and to Luca Chiantini and Jarek Buczyński for related discussions as well. The first author also wishes to express his gratitude to Hajime Kaji and Hiromichi Takagi for useful discussions.
@@ 2.1. Projection from the incidence to the secant variety
(3) $I^{0}=I_{(n)}^{0}:=\left\{(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
where we write $x_{i}=v_{d}(x_{i}^{\prime})$ for $x_{i}^{\prime}\in{\mathbb{P}}^{n}$.
[…]
Taking the first projection $p:I\rightarrow{\mathbb{P}}^{N}$, we have $p(I)=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (see also [28, definition 1.1.3], [29, chapter I, §1]). For any $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, $p^{-1}(a)$ is often called the secant fiber of $a$. Note that $I$ is invariant under permuting factors on $({\mathbb{P}}^{n})^{k}$ from the definition so that both $p$ and $q_{i}$ maps factor through ${\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$. We also have $\dim I={nk+k-1}$ by considering general fibers of $q:I\rightarrow({\mathbb{P}}^{n})^{k}$. For each $1\leq i\leq k$, let
@@
be the composition of $q$ and the projection to the $i$-th factor $({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$.
@@
We have some remarks on the incidence $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$.
@@
    $I^{0}$ can be viewed as a ${\mathbb{P}}^{k-1}$-bundle over a non-empty open subset $U$ of $({\mathbb{P}}^{n})^{k}$, consisting of $k$-tuples of points with the expected spanning dimension, so that $I$ is irreducible. Further, $I^{0}=q^{-1}(U)$, since both are irreducible closed subsets in ${\mathbb{P}}^{N}\times U$ having the same dimension. So, for any $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ and for $x_{i}=v_{d}(x_{i}^{\prime})$, it holds that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ (in other words, there is no $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ such that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1$ but $a\notin\langle\,x_{1},\dots,x_{k}\,\rangle$). Finally, note that the Euclidean closure of $I^{0}$ also coincides with $I$ in this case (see e.g. [22, theorem 3.1.6.1]).
@@
    In the case of $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, it holds $p(I\setminus I^{0})\neq\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In addition, if $k\leq\binom{n+d-1}{n}$, then as setting
@@
    (Alternative incidences for the $k$-th secant variety) In the literature, instead of $(X)^{k}$, other spaces such as the symmetric product $\operatorname{Sym}^{k}(X)$ (e.g. in [9]), the Hilbert scheme of degree $k$ 00-dimensional subschemes $\operatorname{Hilb}_{k}(X)$ (e.g. in [7]), and the Grassmannian $\mathbb{G}(k-1,N)$ (e.g. in [26]) also have been used in the incidence to consider the $k$-th secant variety of $X\subset{\mathbb{P}}^{N}$.
@@
Let us fix an $m$-plane ${\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ and consider the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$, $v_{d}:{\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta}$ with $\beta={m+d\choose m}-1$. We often use the following notation.
@@
    Changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}:=n-m$, we may assume that a given ${\mathbb{P}}^{m}$ in ${\mathbb{P}}^{n}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$. For any $x_{i}^{\prime}\in{\mathbb{P}}^{n}$, say
@@
    then we set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$. Thus, $y_{i}^{\prime}$ gives a point of the $m$-plane if $x_{i}^{\prime}$ is not of the form [0::0:::][0$\dots$:0$\ast$:$\dots$:$\ast$][ 0 : … : 0 : ∗ : … : ∗ ]. By abuse of notations, we denote by $y_{i}^{\prime}$ both corresponding points in ${\mathbb{P}}^{m}$ and in ${\mathbb{P}}^{n}$. Further, as considering linear projections $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ (eliminating the $u$-variables) and $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ (eliminating all the monomials of degree $d$ which involve $u$-variables), we have a natural commuting diagram as $v_{d}\circ\pi_{1}=\pi_{2}\circ v_{d}$
@@
Let us study behaviors of some points in the boundary of $I$, which belong to $I$ but do not belong to $I^{0}$, as follows.
@@
Let $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be any $m$-plane with $m<n$, and consider ${\mathbb{P}}^{\beta}=\langle\,v_{d}(L)\,\rangle\subset{\mathbb{P}}^{N}$. For $v_{d}:L={\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta}$, let us take the incidence variety $I_{(m)}\subset{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ of the same kind as $I=I_{(n)}$ in (3). In this setting, for any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ with $a\in{\mathbb{P}}^{\beta}$, one of the following conditions holds:
@@
Let ${\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ be a linear projection whose center does not intersect with $L$. It induces a natural linear projection ${\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$, and then we have
@@
${\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}},\quad{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\hookrightarrow{\mathbb{P}}^{l_{2}},$
we may regard $\rho_{1}$ as the restriction of a linear projection ${\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$ whose center is a certain linear subvariety $R\subset{\mathbb{P}}^{l_{1}}$.

### v3

@@ 1. Introduction
    (i) of Theorem 1 is stronger than (i) of Theorems 2 and 3, since it claims smoothness for ‘every point’ in the $m$-subsecant locus $\Sigma_{k,d}(m)$ outside $\sigma_{k-1}(v_{d}({\mathbb{P}}^{n}))$. The ‘general point’ condition of (i) in Theorems 2 and 3 cannot be deleted (see Example 27). We also note that, the ranges of $k$ of (i) and (ii) are slightly different between Theorems 1 and 2. If $m=1$ and $d$ is even, then the case $k=\lceil{\binom{m+d}{m}}/({m+1})\rceil=(d+2)/2$ is (ii) of Theorem 1, which is similar to that of Theorem 2; for this $k$, $\sigma_{k}(v_{d}({\mathbb{P}}^{1}))$ is not generically identifiable. However, if $d$ is odd, then the case $k={\binom{m+d}{m}}/({m+1})=(d+1)/2$ belongs to (i) of Theorem 1.
@@
    (i) in Theorems 1, 2, and 3 corresponds to the $k$-identifiable case of a general point $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$. From the viewpoint of the secant fiber in the incidence (4) in §2, it means that the fiber $p^{-1}(a)$ under the projection $p$ consists of only one element up to permuting $x_{i}$’s. On the other hand, (ii) of Theorems 1, 2, and 3 corresponds to the case of generic non-identifiability of the subsecant variety $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ with the situation of $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}=\langle\,v_{d}({\mathbb{P}}^{m})\,\rangle$, except $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$. This non-identifiability of a general $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ occurs if $\dim p^{-1}(a)>0$, or if $\dim p^{-1}(a)=0$ and #$\#p^{-1}(a)\geq 2$ (modulo permutation). For $m=1$, only the former occurs in the case $k=(d+2)/2$. For $m\geq 2$, the former corresponds to the case where $k$ is the ceiling of ${\binom{m+d}{m}}/({m+1})\notin{\mathbb{N}}$ with $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$, to the case where $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is secant defective, or to the case where $\sigma_{k-1}(v_{d}({\mathbb{P}}^{m}))\subsetneq{\mathbb{P}}^{\beta}$ is secant defective (i.e., $(k,d,m)=(8,3,4),(6,4,2),(10,4,3),(15,4,4)$), and the latter corresponds to the case where $k$ is the number ${\binom{m+d}{m}}/({m+1})\in{\mathbb{N}}$ with $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))={\mathbb{P}}^{\beta}$ except $(k,d,m)=(5,3,3),(7,5,2)$, or to the case $(k,d,m)=(9,3,5),(8,4,3),(9,6,2)$.
@@
    $(k,d,m,n)=(3,4,1,2),(4,3,2,3)$ (i.e., (iv) of Theorems 1 and  2), a posteriori, turn out to be the only two exceptional cases which do not follow this trichotomy pattern; in other words, though $k$ belongs to the range of (ii) and the generic non-identifiability of $\sigma_{k}(v_{d}({\mathbb{P}}^{m}))$ holds, $\Sigma_{k,d}(m)$ does not provide non-trivial singular points (see also Remark 24). Indeed, in [FH24] we show that if $(k,d,n)=(3,4,2),(4,3,3)$, then $\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ is a del Pezzo $k$-th secant variety, that is, a $k$-th secant variety of next-to-minimal degree. In this sense, these two cases also belong to a special class with respect to the degrees of higher secant varieties. (For basic definitions and results on such varieties, see [CR], [CK].)
As an application of our main results for $\Sigma_{4,d}(2)$ and $\Sigma_{4,d}(1)$, we obtain the following result on the singularity of the fourth secant variety of any Veronese variety.
@@
    For $d\geq 7$, all points in $\Sigma_{4,d}(1)\setminus\sigma_{3}(v_{d}({\mathbb{P}}^{n}))\subset\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ are smooth. When $d=6$, $\Sigma_{4,d}(1)\subset\operatorname{Sing}(\sigma_{4}(v_{d}({\mathbb{P}}^{n})))$ but $\Sigma_{4,d}(1)\not\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., non-trivial singularity). When $d\leq 5$, $\Sigma_{4,d}(1)\subset\sigma_{3}(v_{d}({\mathbb{P}}^{n}))$ (i.e., trivial singularity).
For a projective variety $X\subset{\mathbb{P}}^{N}$, we denote by $\operatorname{Vertex}(X)$ the set of vertices of $X$. Then $X$ is a cone if and only if $\operatorname{Vertex}(X)\neq\emptyset$.
@@
The smallest case for the singular locus of $\sigma_{4}(v_{d}({\mathbb{P}}^{n}))$ beyond the classical results for $d=2$ or $n=1$ is $(d,n)=(3,2)$, but in this case there is nothing to check, because $\sigma_{4}(v_{3}({\mathbb{P}}^{2}))$ fills up the ambient space ${\mathbb{P}}^{9}$ and then $\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{2})))=\emptyset$. In the case $(d,n)=(3,3)$, by Theorem 5 (i) and (ii), we have
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}^{3})))=\sigma_{3}(v_{3}({\mathbb{P}}^{3})).$
For the case $(d,n)=(3,4)$, if $V$ denotes a $5$-dimensional ${\mathbb{C}}$-vector space with ${\mathbb{P}}^{4}={\mathbb{P}}V$, Theorem 5 tells us that the singular locus of the $4$-th secant variety of $v_{3}({\mathbb{P}}V)$ in ${\mathbb{P}}^{34}$ is precisely the locus of cubic hypersurfaces in $5$ variables which are cones with the vertex dimension $\geq 1$ as follows:
$\operatorname{Sing}(\sigma_{4}(v_{3}({\mathbb{P}}V)))=\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\Sigma_{4,3}(2;{{\mathbb{P}}V})\\ =\sigma_{3}(v_{3}({\mathbb{P}}V))\cup\big{\{}\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\sigma_{4}(v_{3}({\mathbb{P}}^{2}))\big{\}}=\bigcup_{{\mathbb{P}}^{2}\subset{\mathbb{P}}V}\langle\,v_{3}({\mathbb{P}}^{2})\,\rangle\\ =\big{\{}f\in{\mathbb{P}}S^{3}V~{}|~{}\textrm{the cubic hypersurface $X\subset{\mathbb{P}}V$ defined by $f$ is a cone with $\dim\operatorname{Vertex}(X)\geq 1$}\big{\}},$
@@
Such a simple description of the singular locus can be attained in a few more cases (see Corollary 32 for details).
The paper is structured as follows. In §2, as preparation, we first recall some preliminaries on $k$-th secant varieties and corresponding incidences. Then, using projective techniques, such as Terracini’s lemma, the trisecant lemma, descriptions of embedding tangent spaces, and tangential projections, we reveal several geometric properties of $m$-subsecant varieties in higher secant varieties of Veronese varieties, which are crucial for the proof of main theorems. In §3, as an illustration of the whole picture and our main ideas, we treat the case $m=1$ and prove Theorem 1. In §4 we deal with the general case (i.e., $m\geq 2$) and prove Theorem 2 and Theorem 3 as generalizing the ideas used in the previous section. We would like to remark that this can be done because the dimension theorem [AH95] and the generic identifiability question [COV17, GM17] were settled for the case of Veronese varieties. In §5 we focus on the singular locus of the $4$-th secant variety and prove Theorem 5 by dividing the case into two parts; ‘full-secant locus points (i.e., $m=3$)’ treated in Theorem 29 via Young flattening and conormal space computation and ‘the subsecant locus’ by Corollary 30. Finally, we make some generalizations and remarks on the material for further developments in §6.
@@ Acknowledgements
The authors would like to give thank to Korea Institute for Advanced Study (KIAS) for inviting them and giving a chance to start the project. We are grateful to Giorgio Ottaviani for a helpful comment on Example 6 and to Luca Chiantini and Jarek Buczyński for related discussions as well. The first author also wishes to express his gratitude to Hajime Kaji and Hiromichi Takagi for useful discussions. Finally, we would like to thank the anonymous referees very much for their careful reading, pointing out some error in the first version, and many suggestions which help us to improve the final exposition of the paper.
@@ 2.1. Projection from the incidence to the secant variety
(4) $I^{0}=I_{(n)}^{0}=\left\{(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\mid a\in\langle\,x_{1},\dots,x_{k}\,\rangle\text{ and }\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1\right\},$
where we write $x_{i}=v_{d}(x_{i}^{\prime})$ for $x_{i}^{\prime}\in{\mathbb{P}}^{n}$. Taking the first projection $p:I\rightarrow{\mathbb{P}}^{N}$, we have $p(I)=\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$ (see also [Rus16, Definition 1.1.3], [Zak93, Chapter I, §1, Chapter V]). For any $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, $p^{-1}(a)$ is often called the secant fiber of $a$. Note that $I$ is invariant under permuting factors on $({\mathbb{P}}^{n})^{k}$ from the definition so that both $p$ and $q_{i}$ maps factor through ${\mathbb{P}}^{N}\times\operatorname{Sym}^{k}({\mathbb{P}}^{n})$.
We also have $\dim I={nk+k-1}$ by considering general fibers of $q:I\rightarrow({\mathbb{P}}^{n})^{k}$. For each $1\leq i\leq k$, let
@@
be the composition of $q$ and the projection to the $i$-th factor $({\mathbb{P}}^{n})^{k}\rightarrow{\mathbb{P}}^{n}$. Then the following commutative diagram is obtained:
${I}$${{\mathbb{P}}^{N}}$${({\mathbb{P}}^{n})^{k}}$${{\mathbb{P}}^{n}.}$$\scriptstyle{p}$$\scriptstyle{q}$$\scriptstyle{q_{i}}$
@@
We have some remarks on the incidence variety $I\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}$.
@@
    $I^{0}$ can be viewed as a ${\mathbb{P}}^{k-1}$-bundle over a non-empty open subset $U$ of $({\mathbb{P}}^{n})^{k}$, consisting of $k$-tuples of points with the expected spanning dimension, so that $I$ is irreducible. Further, for $q^{-1}(U)=I\cap({\mathbb{P}}^{n}\times U)$, we have $I^{0}=q^{-1}(U)$ since both are irreducible closed subsets in ${\mathbb{P}}^{N}\times U$ having the same dimension. So, for any $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ and for $x_{i}=v_{d}(x_{i}^{\prime})$, it holds that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle<k-1$ (in other words, there is no $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in I\setminus I^{0}$ such that $\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1$ but $a\notin\langle\,x_{1},\dots,x_{k}\,\rangle$). Finally, note that the Euclidean closure of $I^{0}$ also coincides with $I$ in this case (see e.g. [Land17, Theorem 3.1.6.1]).
@@
    In the case of $\dim\sigma_{k}(v_{d}({\mathbb{P}}^{n}))={nk+k-1}$, it holds $p(I\setminus I^{0})\neq\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$, and then $p^{-1}(a)\subset I^{0}$ for general $a\in\sigma_{k}(v_{d}({\mathbb{P}}^{n}))$. In addition, if $k\leq\binom{n+d-1}{n}$, then as setting
@@
    (Alternative incidences for the $k$-th secant variety) In the literature, instead of $(X)^{k}$, other spaces such as the symmetric product $\operatorname{Sym}^{k}(X)$ (e.g. in [COV17]), the Hilbert scheme of degree $k$ 00-dimensional subschemes $\operatorname{Hilb}_{k}(X)$ (e.g. in [BB14]), and the Grassmannian $\mathbb{G}(k-1,N)$ (e.g. in [Mel06]) have also been used in the incidence to consider the $k$-th secant variety of $X\subset{\mathbb{P}}^{N}$.
@@
Let us fix an $m$-plane $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ and consider the $d$-uple Veronese embedding of ${\mathbb{P}}^{m}$, $v_{d}:L={\mathbb{P}}^{m}\rightarrow{\mathbb{P}}^{\beta}$ with $\beta={m+d\choose m}-1$. We often use the following notation.
@@
    Let $\hat{L}\subset{\mathbb{P}}^{n}$ be any $(n-m-1)$-plane not intersecting $L$. Changing homogeneous coordinates $t_{0},t_{1},\dots,t_{m},u_{1},u_{2},\dots,u_{m^{\prime}}$ on ${\mathbb{P}}^{n}$ with $m^{\prime}=n-m$, we may assume that $L\subset{\mathbb{P}}^{n}$ is the zero set of $u_{1},\dots,u_{m^{\prime}}$ and $\hat{L}\subset{\mathbb{P}}^{n}$ is the zero set of $t_{0},\dots,t_{m}$. For any $x_{i}^{\prime}\in{\mathbb{P}}^{n}$, say
@@
    then we set $y_{i}^{\prime}=[x_{i,0}^{\prime}:\dots:x_{i,m}^{\prime}:0:\dots:0]$. Thus, $y_{i}^{\prime}$ gives a point of the $m$-plane if $x_{i}^{\prime}$ is not of the form [0::0:::][0$\dots$:0$\ast$:$\dots$:$\ast$][ 0 : … : 0 : ∗ : … : ∗ ]. By abuse of notations, we denote by $y_{i}^{\prime}$ both corresponding points in ${\mathbb{P}}^{m}$ and in ${\mathbb{P}}^{n}$. Further, as considering linear projections $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ from the center $\hat{L}$ (eliminating the $u$-variables), and $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ (eliminating all the monomials of degree $d$ which involve $u$-variables), we have a natural commuting diagram as $v_{d}\circ\pi_{1}=\pi_{2}\circ v_{d}$
@@
Let us study the behavior of some points in the boundary of $I$, which belong to $I$ but do not belong to $I^{0}$, as follows.
@@
Let $L={\mathbb{P}}^{m}\subset{\mathbb{P}}^{n}$ be any $m$-plane with $m<n$, and consider ${\mathbb{P}}^{\beta}=\langle\,v_{d}(L)\,\rangle\subset{\mathbb{P}}^{N}$. For $v_{d}:L={\mathbb{P}}^{m}\hookrightarrow{\mathbb{P}}^{\beta}$, let us take $I_{(m)}\subset{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ to be the closure of the set of points $(a,{x_{1}^{\prime},\dots,x_{k}^{\prime}})\in{\mathbb{P}}^{\beta}\times({\mathbb{P}}^{m})^{k}$ such that $a\in\langle\,x_{1},\dots,x_{k}\,\rangle$ and $\dim\langle\,x_{1},\dots,x_{k}\,\rangle=k-1$ (i.e., the incidence variety of the same kind as $I=I_{(n)}$ in (4)). In this setting, for any $(a,x_{1}^{\prime},\dots,x_{k}^{\prime})\in I\setminus I^{0}$ with $a\in{\mathbb{P}}^{\beta}$, one of the following conditions holds:
@@
Let $\left\{W_{j}\right\}$ be the irreducible components of $I\setminus I^{0}$. For each $j$, we take
(6) $(a_{j},x_{j,1}^{\prime},\dots,x_{j,k}^{\prime})\in W_{j}\subset{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}.$
Let $\hat{L}\subset{\mathbb{P}}^{n}$ be a general $(n-m-1)$-plane such that all $x_{j,i}^{\prime}\notin\hat{L}$ and $L\cap\hat{L}=\emptyset$. For the linear projection $\pi_{1}:{\mathbb{P}}^{n}\dashrightarrow{\mathbb{P}}^{m}$ from the center $\hat{L}$, we have a natural linear projection $\pi_{2}:{\mathbb{P}}^{N}\dashrightarrow{\mathbb{P}}^{\beta}$ as in Remark 8 with the diagram (5), and then define the projections
@@
$Seg_{1}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{n})^{k}\hookrightarrow{\mathbb{P}}^{l_{1}},\quad Seg_{2}:{\mathbb{P}}^{N}\times({\mathbb{P}}^{m})^{k}\hookrightarrow{\mathbb{P}}^{l_{2}},$
we may regard $\rho_{1}$ as the restriction of a linear projection $\pi_{R}:{\mathbb{P}}^{l_{1}}\dashrightarrow{\mathbb{P}}^{l_{2}}$ whose center $R$ is a certain linear subvariety of ${\mathbb{P}}^{l_{1}}$. Let ~$\tilde{{\mathbb{P}}}^{l_{1}}\subset{\mathbb{P}}^{l_{1}}\times{\mathbb{P}}^{l_{2}}$ be the graph of $\pi_{R}$, which coincides with the blowing-up of ${\mathbb{P}}^{l_{1}}$ with respect to $R$. Let $r_{1}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{1}}$ and $r_{2}:\tilde{{\mathbb{P}}}^{l_{1}}\rightarrow{\mathbb{P}}^{l_{2}}$ be projections. Then we have the following commutative diagram: